# Computing High-dimensional Confidence Sets for Arbitrary Distributions

### v1

@@ 1.3 Application to Conformal Prediction
Theorem 1.5 (Conformal Prediction with Approximate Restricted Volume Optimality).
We have an algorithm for conformal prediction over examples from $\mathcal{Y}=\mathbb{R}^{d}$, that achieves approximate restricted volume optimality444an approximate form of restricted volume optimality defined by Gao et al. (2025). with respect to the set $\mathcal{C}$ of Euclidean balls. That is, we have an algorithm that, given training examples $Y_{1},\dots,Y_{n}\in\mathcal{Y}$, miscoverage rate $10<\alpha<$ 1, and coverage slack factor $0\leq\gamma\leq 1$, outputs a set $\widehat{C}$ (not necessarily in $\mathcal{C}$), such that for an unknown test example $Y_{n+1}\in\mathcal{Y}$,
@@ 1.4 Prior and Related Algorithmic Approaches
Coresets are a powerful algorithmic primitive for geometric problems involving an input set of data points (Agarwal et al., 2004). A coreset for a problem is a small summation of the data (e.g., a subset of the points) such that solving the problem on the summation gives an approximate solution to the whole instance. Badiou, Har-Peled and Indyk proved a surprising result about coresets for the problem of minimum volume enclosing ball for a set of points – they showed that there is a coreset of size $1/\varepsilon$ (that is independent of the dimension) whose minimum volume enclosing ball approximates the minimum volume ball enclosing all the points up to a $(1+\varepsilon)$ factor in radius (Badoiu et al., 2002). [^5] The coreset size of $\lceil 1/\varepsilon\rceil$ is also known to be tight (Agarwal et al., 2004). Badoiu et al. (2002) show how to find a ball enclosing $\delta-\gamma$ fraction of the given points in polynomial time whose radius is within a factor $\Gamma=\Big{(}1+O_{\gamma}(\frac{(\log\log d)^{2}}{\log d})\Big{)}$ (see Theorem 4.2 of Badoiu et al., 2002).  (See also Ding, 2020; Agarwal et al., 2004). This corresponds to an approximation factor of $\Gamma^{d}=\exp\big{(}\widetilde{O}(d/\log d)\big{)}$ for the volume. We are not aware of any other work that get a better competitive ratio $\Gamma$ for balls.
@@ 2.1 Proper Learning for Non-Worst-Case Instances
To tackle (a) above, our main observation is that we can approximate the center $c^{\star}$ of the minimum volume ball $B^{\star}$ containing $Y^{\star}$ by using the mean $\mu^{\star}$ of $Y^{\star}$. It is not in general true that $\mu^{\star}$ should be near $c^{\star}$. However, if $\mu^{\star}$ is far from $c^{\star}$, most points in $Y^{\star}$ must be closer to $\mu^{\star}$ than they are to $c^{\star}$. This is illustrated in Figure 1.
Figure 1: The figure shows the ball $B^{\star}$ with center $c^{\star}$, but the mean of the points $\mu^{\star}$ is far from the center $c^{\star}$. In this case, most of the points in $B^{\star}$ are actually closer to $\mu^{\star}$ than to $c^{\star}$.
@@
Let $Y$ be a set of points in $\mathbb{R}^{d}$ with mean $\mu_{Y}$ and covariance matrix $\Sigma$, where the largest eigenvalue of $\Sigma$ satisfies $\lambda_{\max}(\Sigma)\leq\sigma^{2}$, and such that $Y$ is contained in a ball $B(c,R)$. Then for any $t\geq 1$, there exists a subset $\widehat{Y}\subseteq Y$, such that $|\widehat{Y}|\geq\left(1-\frac{1}{t^{2}}\right)|Y|$, and for every point $y\in\widehat{Y}$,
$\|y-\mu_{Y}\|\leq\|y-c\|+2t\sigma.$
@@
We assume without loss of generality that $c$ is the origin, and therefore $y-c=y$ for all $y\in\mathbb{R}^{d}$. Thus, we aim to show that there exists a subset $\widehat{Y}\subseteq Y$, such that $|\widehat{Y}|\geq\left(1-\frac{1}{t^{2}}\right)|Y|$, and for every point $y\in\widehat{Y}$,
$\|y-\mu_{Y}\|\leq\|y\|+2t\sigma.$
We consider two cases: (1) $\|\mu_{Y}\|\geq 2t\sigma$; and (2) $\|\mu_{Y}\|<2t\sigma$.
Case 1: Suppose $\|\mu_{Y}\|\geq 2t\sigma.$
We start by bounding the number of points that deviate significantly from the mean using Chebyshev’s inequality. Let $v_{\mu}=\frac{\mu_{Y}}{||\mu_{Y}||}$ be the unit vector in the direction of $\mu_{Y}$. We have that, for $y$ sampled uniformly from $Y$, i.e. $y\sim\mathrm{Unif}(Y)$,
$\mathbf{P}\left[|\langle y-\mu_{Y},v_{\mu}\rangle|\geq t\sigma\right]\leq\frac{1}{t^{2}}.$
This means that there exists a subset $\widehat{Y}\subseteq Y$, such that $|\widehat{Y}|\geq\left(1-\frac{1}{t^{2}}\right)|Y|$, and for $y\in\widehat{Y}$ we have $|\langle y-\mu_{Y},v_{\mu}\rangle|\leq t\sigma.$
Now, we decompose each $y\in\widehat{Y}$ into the portion that is parallel to $\mu_{Y}$,
$y^{\parallel}=\langle y,v_{\mu}\rangle v_{\mu},$
and the portion that is orthogonal to $\mu_{Y}$,
$y^{\perp}=y-y^{\parallel}.$
Then for the parallel portion, since $y\in\widehat{Y}$, we have that
$\|y^{\parallel}-\mu_{Y}\|_{2}$ $=\|\langle y,v_{\mu}\rangle v_{\mu}-\mu_{Y}\|_{2}=\left|\langle y-\mu_{Y},v_{\mu}\rangle\right|\leq t\sigma.$
By the triangle inequality, we have
$\|y^{\parallel}\|_{2}\geq\|\mu_{Y}\|_{2}-\|y^{\parallel}-\mu_{Y}\|_{2}\geq 2t\sigma-t\sigma\geq t\sigma\geq\|y^{\parallel}-\mu_{Y}\|_{2}.$
Finally, this concludes that for every point $y\in\widehat{Y}$,
$\|y-\mu_{Y}\|_{2}^{2}=\|y^{\parallel}-\mu_{Y}\|_{2}^{2}+\|y^{\perp}\|_{2}^{2}\leq\|y^{\parallel}\|_{2}^{2}+\|y^{\perp}\|_{2}^{2}=\|y\|_{2}^{2},$
which satisfies the required condition.
Case 2: Suppose $\|\mu_{Y}\|\leq 2t\sigma.$
We set $\widehat{Y}$ to include all of the points in $Y$, i.e. $\widehat{Y}=Y$. To bound the distance from each point in $Y$ to $\|\mu_{Y}\|$, we use the triangle inequality. Specifically, for every $y\in\widehat{Y}$, we have
$\|y-\mu_{Y}\|\leq\|y\|+\|\mu_{Y}\|\leq\|y\|+2t\sigma,$
which completes the proof. ∎
@@ 2.2 Reducing Worst-Case Improper Learning to Non-Worst-Case Proper Learning
Suppose we are given a set of points $Y\subseteq\mathbb{R}^{d},|Y|=n$, and $0\leq\delta\leq 1,0\leq\gamma\leq 1$, such that there exists a subset $Y^{\star}\subseteq Y$, $|Y^{\star}|\geq\delta|Y|$, that is contained in an unknown ball $B^{\star}=B(c^{\star},R^{\star})$. Then we can find an ellipsoid $\widehat{E}$ such that $|\widehat{E}\cap Y|\geq\delta(1-\gamma)|Y^{\star}|$, and
$\mathrm{vol}(\widehat{E})^{1/d}\leq\mathrm{vol}(B^{\star})^{1/d}\cdot\left(1+d^{-1/3+o(1)}\cdot O\left(\frac{1}{\delta}+\frac{1}{\sqrt{\gamma\delta}}\right)\right).$
Intuitively, the case where our non-worst-case proper learning algorithm fails, is when the points have most of their variance concentrated in only a few directions, see for example Figure 2. This makes it hard to estimate the position of the center accurately in the large variance directions. However, this also means that the variance of the points in the other directions must be low, so it is easier to estimate the position of the center in these directions. Applying the correct linear transformation to the points essentially allows us to “reweigh” the accuracy that we are aiming for across the different directions, and trade off error in the high variance directions for accuracy in the low variance directions.
@@
Our strategy is to find a good preconditioner $\widehat{M}^{-1/2}$, for the points in $Y$, such that the transformed points have small variance. (Writing the preconditioner as $\widehat{M}^{-1/2}$ is convenient, as we can then argue about the distortion in terms of the eigenvalues of $\widehat{M}$, which we will choose to be positive semidefinite.) Then, we can apply our procedure to find an approximately minimum volume ball on the transformed points, which will return a strong approximation since the variance of the points is controlled (via Lemma 2.1). Transforming this ball back into the original space will result in a confidence region that is shaped as an ellipsoid.
@@
    $\widehat{M}$ is positive semidefinite, and therefore $\widehat{M}^{-1/2}$ is “shrinking” (non-expanding) in all directions. This ensures that there exists a ball in the transformed space that contains $Y^{\star}$ and has volume at most that of $B^{\star}$ in the original space.
@@
Properties (i) and (ii) allow us to reduce to the non-worst-case setting, and property (iii) ensures that this translation does not lose too much in volume. We will set $\widehat{\tau}=d^{1/6}$ to trade-off the different losses from (ii) and (iii) and obtain the final guarantee. The full proof of this theorem can be found in Section 4.
@@ 3 Proper Learning of Euclidean Balls
We use this result as a subroutine in our main result for improper learning. In this section we provide some additional results that extend our result for proper learning to other settings. In particular, Theorem 3.1 gives a proper learning algorithm for fully worst-case instances, and achieves an approximation ratio of $\left(1+\widetilde{O}_{\gamma,\delta}\left(\frac{1}{\sqrt{\log d}}\right)\right)$ in radius, which is slightly weaker than the best known of $\left(1+\widetilde{O}_{\gamma,\delta}\left(\frac{1}{\log d}\right)\right)$ due to Badoiu et al. (2002). While we do achieve a slightly weaker approximation guarantee in this setting, our techniques are significantly different than the coreset based techniques of the prior work, and extend well to non-worst-case settings, as evidenced by our improper learning result. We also provide Theorem 1.4, which shows a way to achieve a better approximation ratio, assuming non-worst-case properties only on the inlier points.
@@
$\operatorname*{\mathrm{vol}}(\widehat{B})^{1/d}\leq\operatorname*{\mathrm{vol}}(B^{\star})^{1/d}\left(1+O\left(\sqrt{\frac{\log\log d}{\log d}}\left(\frac{1}{\sqrt{\gamma}}+\sqrt{(1-\delta)/\delta}\right)\right)\right)$
@@
$\operatorname*{\mathrm{vol}}(\widehat{B})^{1/d}\leq\operatorname*{\mathrm{vol}}(B^{\star})^{1/d}\left(1+O\left(\sqrt{\frac{\log\log d}{\log d}}\left(\frac{1}{\sqrt{\gamma}}+\sqrt{(1-\delta)/\delta}\right)\right)\right)$
@@
Our approach starts with the following structural observation about the true high-density ball $B^{\star}$. It is not in general true that the mean $\mu^{\star}$ of the points $Y^{\star}=B^{\star}\cap Y$ should be near the center $c^{\star}$ of $B^{\star}$. However, it turns out that if the mean is far from the center, taking a ball centered at the mean of $B^{\star}$ must still capture most of the points in $Y^{\star}$.
This means that by taking a ball centered at $\mu^{\star}$, either (1) $\mu^{\star}$ is close to $c^{\star}$, and a slightly larger radius than $B^{\star}$ captures all of the points in $B^{\star}$, or (2) $\mu^{\star}$ is far and taking a ball of equal radius centered at $\mu^{\star}$ captures most of the points in $B^{\star}$. Here, the notions of close and far depend on the variance of the points in $Y^{\star}$. Thus our problem reduces to the following. Given only the full set of points $Y$, we must estimate the mean of $Y^{\star}$, while also controlling the variance of the points that we are taking the mean of, to minimize the extra radius we pay for in case (1).
@@
Suppose there exists a subset of points $Y^{\star}\subseteq Y^{\prime}$ are contained in a smaller ball $B^{\star}=B(c^{\star},R^{\star})$, where $|Y^{\star}|\geq\delta|Y^{\prime}|$ and $R^{\prime}\leq 2R^{\star}$.
Then there exists a list $C$ of size at most $\left(\frac{3R^{\prime}}{\tau}\right)^{q}$, such that there exists a point $\widehat{c}\in C$ such that for
$\widehat{R}\leq R^{\star}+2\sigma\left(\frac{2}{\sqrt{\gamma}}+\sqrt{2(1-\delta)/\delta}\right)+2\tau,$
the ball $\widehat{B}=B(\widehat{c},\widehat{R})$ contains a subset $\widehat{Y}\subseteq Y^{\prime}$ such that $|\widehat{Y}|\geq(1-\gamma)|Y^{\star}|$.
@@
We start by considering $\Sigma_{Y^{\prime}}$, the covariance matrix of $Y^{\prime}$ and partition its eigenvalues into two groups. The directions associated with the largest $q$ eigenvalues of $\Sigma_{Y^{\prime}}$ are referred to as the “high variance directions”; while the directions corresponding to the remaining eigenvalues are the “low variance directions”. Let $\Pi_{\mathrm{high}}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{q}$ denote the projection onto the “high variance” eigenspace and $\Pi_{\mathrm{low}}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d-q}$ denote the projection onto the “low variance” eigenspace. We then estimate the center $\widehat{c}$ of the ball $\widehat{B}$.
First, we estimate the location of $\widehat{c}$ in the high variance directions. Let $Y^{\prime}_{\mathrm{high}}=\Pi_{\mathrm{high}}Y^{\prime}$ denote the projection of points in $Y^{\prime}$ onto the high variance directions. Note that the set $Y^{\prime}_{\mathrm{high}}$ lies within the $q$-dimensional ball $B^{\prime}_{\mathrm{high}}=\Pi_{\mathrm{high}}B^{\prime}$, which is centered at the origin and has a radius of at most $R^{\prime}$. Next, consider the projection of $Y^{*}$ onto the high variance directions, denoted by $Y^{*}_{\mathrm{high}}=\Pi_{\mathrm{high}}Y^{*}$. Then, this set $Y^{*}_{\mathrm{high}}$ is contained in the ball $B^{*}_{\mathrm{high}}=\Pi_{\mathrm{high}}B^{*}$ centered at $\Pi_{\mathrm{high}}c^{*}$ with radius at most $R^{\star}$. By Lemma 3.5, there exists a list $L\subset\mathbb{R}^{q}$ with size $|L|\leq\left(\frac{2+\tau/R^{\prime}}{\tau/R^{\prime}}\right)^{q}\leq\left(\frac{3R^{\prime}}{\tau}\right)^{q}$ such that there is a point ^$\widehat{\ell}\in L$ such that for every point $y\in Y^{\star}$,

### v2

@@ 1.3 Application to Conformal Prediction
Theorem 1.5 (Conformal Prediction with Approximate Volume Optimality).
We have an algorithm for conformal prediction over examples from $\mathcal{Y}=\mathbb{R}^{d}$, that achieves approximate volume optimality with respect to the set $\mathcal{C}$ of Euclidean balls. [^4] This is an approximate form of restricted volume optimality defined by Gao et al. (2025). That is, we have an algorithm that, given training examples $Y_{1},\dots,Y_{n}\in\mathcal{Y}$, miscoverage rate $10<\alpha<$ 1, and coverage slack factor $0\leq\gamma\leq 1$, outputs a set $\widehat{C}$ (not necessarily in $\mathcal{C}$), such that for an unknown test example $Y_{n+1}\in\mathcal{Y}$,
@@
    $\mathrm{vol}(\widehat{C})^{1/d}\leq\left(1+O_{\gamma,\delta}\big{(}d^{-1/2+o(1)}\big{)}\right)\mathrm{vol}(C^{\star})^{1/d},$
@@ 1.4 Prior and Related Algorithmic Approaches
Coresets are a powerful algorithmic primitive for geometric problems involving an input set of data points (Agarwal et al., 2004). A coreset for a problem is a small summation of the data (e.g., a subset of the points) such that solving the problem on the summation gives an approximate solution to the whole instance. Badiou, Har-Peled and Indyk proved a surprising result about coresets for the problem of minimum volume enclosing ball for a set of points – they showed that there is a coreset of size $1/\varepsilon$ (that is independent of the dimension) whose minimum volume enclosing ball approximates the minimum volume ball enclosing all the points up to a $(1+\varepsilon)$ factor in radius (Badoiu et al., 2002). [^5] The coreset size of $\lceil 1/\varepsilon\rceil$ is also known to be tight (Agarwal et al., 2004). Badoiu et al. (2002) show how to find a ball enclosing $\delta-\gamma$ fraction of the given points in polynomial time whose radius is within a factor $\Gamma=\Big{(}1+O_{\gamma}(\frac{(\log\log d)^{2}}{\log d})\Big{)}$ (see Theorem 4.2 of Badoiu et al., 2002). The subsequent work of Ding (2020) also uses coresets and seems to obtain a guarantee of $\Gamma=\Big{(}1+O_{\gamma}(\frac{1}{\log d})\Big{)}$ (see (24) following Corollary 11 in (Ding, 2020)), though this is not stated explicitly. This corresponds to an approximation factor of $\Gamma^{d}=\exp\big{(}\widetilde{O}(d/\log d)\big{)}$ for the volume. We are not aware of any other work that get a better competitive ratio $\Gamma$ for balls.
@@ 2.1 Proper Learning for Non-Worst-Case Instances
To tackle (a) above, our main observation is that we can approximate the center $c^{\star}$ of the minimum volume ball $B^{\star}$ containing $Y^{\star}$ by using the mean $\mu^{\star}$ of $Y^{\star}$. It is not in general true that $\mu^{\star}$ should be near $c^{\star}$. However, most points in $Y^{\star}$ must be approximately as close to $\mu^{\star}$ as they are to $c^{\star}$. This is illustrated in Figure 1.
Figure 1: The figure shows the points $Y$ (in red) with mean $\mu$, that are contained in ball $B$ with center $c$ and radius $R$. It is not necessarily the case that $\mu$ is near $c$, as $B$ can be defined by just a few points. However, Chebyshev’s inequality tells us that most of the points $Y$ (depicted here as the points in the shaded region) are within a few standard deviations ($\sigma$) of $\mu$ in the $\mu-c$ direction. This allows us to bound the distance of these points to $\mu$ as $\leq\sqrt{R^{2}+(t\sigma)^{2}}$.
@@
Let $Y$ be a set of points in $\mathbb{R}^{d}$ that are contained in a ball $B(c,R)$. Let $\mu_{Y}$ be the empirical mean of $Y$, and $\Sigma_{Y}$ be the empirical covariance, where the largest eigenvalue of $\Sigma_{Y}$ satisifies $\lambda_{\mathrm{max}}(\Sigma_{Y})\leq\sigma^{2}$. Let $\widehat{\mu}$ be an arbitrary point such that $||\mu_{Y}-\widehat{\mu}||_{2}\leq\tau$. Then for any $t\geq 1$, there exists a subset $\widehat{Y}$ such that $|\widehat{Y}|\geq\left(1-\frac{1}{t^{2}}\right)|Y|$, and for every point $y\in\widehat{Y}$,
$||y-\widehat{\mu}||^{2}_{2}\leq||y-c||_{2}^{2}+t^{2}(\sigma^{2}+\tau^{2}).$
@@
As a warm-up, we can bound the number of points that deviate significantly from $\mu_{Y}$ using Chebyshev’s inequality. Let $v_{\mu}=\frac{\mu_{Y}-c}{||\mu_{Y}-c||}$ be the unit vector in the direction of $\mu_{Y}-c$. We have that
$\mathbf{P}_{y\sim\mathrm{Unif}(Y)}\left[|\langle y-\mu_{Y},v_{\mu}\rangle|\geq t\sigma\right]\leq\frac{1}{t^{2}}.$
This means that there exists a subset $\widehat{Y}\subseteq Y$ such that $|\widehat{Y}|\geq\left(1-\frac{1}{t^{2}}\right)|Y|$, and for all $y\in\widehat{Y}$ we have $|\langle y-\mu_{Y},v_{\mu}\rangle|\leq t\sigma$. For each $y\in\widehat{Y}$, we can decompose $y-\mu_{Y}$ into the portion that is parallel to $v_{\mu}$: $\overline{(y-\mu_{Y})}=\langle y-\mu_{Y},v_{\mu}\rangle v_{\mu}$, and the portion that is perpendicular to $v_{\mu}$: $(y-\mu_{Y})^{\perp}=(y-\mu_{Y})-\overline{(y-\mu_{Y})}$. Let $\overline{(y-c)}$ be the portion of $y-c$ that is parallel to $v_{\mu}$: $\overline{(y-c)}=\langle y-c,v_{\mu}\rangle v_{\mu}$. By the choice of $v_{\mu}$, we have that $\overline{(y-c)}=\overline{(y-\mu_{Y})}$. Thus,
$||y-\mu_{Y}||^{2}_{2}=||(y-\mu_{Y})^{\perp}||_{2}^{2}+||\overline{(y-\mu_{Y})}||^{2}_{2}=||(y-c)^{\perp}||_{2}^{2}+||\overline{(y-\mu_{Y})}||^{2}_{2}\leq||y-c||_{2}^{2}+(t\sigma)^{2}.$
Instead of bounding the number of points that deviate significantly from $\mu_{Y}$, we would like to bound the number of points that deviate significantly from $\widehat{\mu}$. We observe that a similar argument goes through. We have that
$\mathbf{E}_{y\sim\mathrm{Unif}(Y)}\left[(y-\widehat{\mu})(y-\widehat{\mu})^{\top}\right]$ $=\mathbf{E}_{y\sim\mathrm{Unif}(Y)}\left[(y-\mu_{Y}+\mu_{Y}-\widehat{\mu})(y-\mu_{Y}+\mu_{Y}-\widehat{\mu})^{\top}\right]$
$=\mathbf{E}_{y\sim\mathrm{Unif}(Y)}\left[(y-\mu_{Y})(y-\mu_{Y})^{\top}\right]+\mathbf{E}_{y\sim\mathrm{Unif}(Y)}\left[(y-\mu_{Y})\right](\mu_{Y}-\widehat{\mu})^{\top}$
$\qquad\qquad+(\mu_{Y}-\widehat{\mu})\mathbf{E}_{y\sim\mathrm{Unif}(Y)}\left[(y-\mu_{Y})^{\top}\right]+(\mu_{Y}-\widehat{\mu})(\mu_{Y}-\widehat{\mu})^{\top}$
$=\Sigma_{Y}+(\mu_{Y}-\widehat{\mu})(\mu_{Y}-\widehat{\mu})^{\top}$
$\preccurlyeq\Sigma_{Y}+\tau^{2}I.$ (7)
Let $v_{\widehat{\mu}}=\frac{\widehat{\mu}-c}{||\widehat{\mu}-c||}$ be a unit vector in the direction of $\widehat{\mu}-c$. By (7) we have that
$\mathbf{P}_{y\sim\mathrm{Unif}(Y)}\left[|\langle y-\widehat{\mu},v_{\widehat{\mu}}\rangle|\geq t\sqrt{\sigma^{2}+\tau^{2}}\right]\leq\frac{1}{t^{2}}.$
Thus there exists a subset $\widehat{Y}\subseteq Y$ such that $|\widehat{Y}|\geq\left(1-\frac{1}{t^{2}}\right)|Y|$, and for all $y\in\widehat{Y}$ we have $|\langle y-\widehat{\mu},v_{\widehat{\mu}}\rangle|\leq t\sqrt{\sigma^{2}+\tau^{2}}$. For each $y\in\widehat{Y}$, we can decompose $y-\widehat{\mu}$ into the portion that is parallel to $v_{\widehat{\mu}}$: $\overline{(y-\widehat{\mu})}=\langle y-\widehat{\mu},v_{\widehat{\mu}}\rangle v_{\widehat{\mu}}$, and the portion that is perpendicular to $v_{\widehat{\mu}}$: $(y-\widehat{\mu})^{\perp}=(y-\widehat{\mu})-\overline{(y-\widehat{\mu})}$. Let $\overline{(y-c)}$ be the portion of $y-c$ that is parallel to $v_{\widehat{\mu}}$: $\overline{(y-c)}=\langle y-c,v_{\widehat{\mu}}\rangle v_{\widehat{\mu}}$. By the choice of $v_{\widehat{\mu}}$, we have that $\overline{(y-c)}=\overline{(y-\widehat{\mu})}$. Thus we have
$||y-\widehat{\mu}||^{2}_{2}=||(y-\widehat{\mu})^{\perp}||_{2}^{2}+||\overline{(y-\widehat{\mu})}||^{2}_{2}=||(y-c)^{\perp}||_{2}^{2}+||\overline{(y-\widehat{\mu})}||^{2}_{2}\leq||y-c||_{2}^{2}+t^{2}(\sigma^{2}+\tau^{2}).$
@@ 2.2 Reducing Worst-Case Improper Learning to Non-Worst-Case Proper Learning
Suppose we are given a set of points $Y\subseteq\mathbb{R}^{d},|Y|=n$, and $0\leq\delta\leq 1,0\leq\gamma\leq 1$, such that there exists a subset $Y^{\star}\subseteq Y$, $|Y^{\star}|\geq\delta|Y|$, that is contained in an unknown ball $B^{\star}=B(c^{\star},R^{\star})$. Then we can find an ellipsoid $\widehat{E}$ such that $|\widehat{E}\cap Y|\geq\delta(1-\gamma)|Y|$, and
$\mathrm{vol}(\widehat{E})^{1/d}\leq\mathrm{vol}(B^{\star})^{1/d}\cdot\left(1+O\left(d^{-1/2+o(1)}/\gamma\delta\right)\right).$
Intuitively, the case where our non-worst-case proper learning algorithm fails is when the points have most of their variance concentrated in only a few directions, see for example Figure 2. This makes it hard to estimate the position of the center accurately in the large variance directions. However, this also means that the variance of the points in the other directions must be low, so it is easier to estimate the position of the center in these directions. Applying the correct linear transformation to the points essentially allows us to “reweigh” the accuracy that we are aiming for across the different directions, and trade off error in the high variance directions for accuracy in the low variance directions.
@@
Our strategy is to find a good preconditioner $\widehat{M}^{-1/2}$, for the points in $Y$, such that the transformed points have small variance. (Writing the preconditioner as $\widehat{M}^{-1/2}$ is convenient, as we can then argue about the distortion in terms of the eigenvalues of $\widehat{M}$, which we will choose to be positive definite.) Then, we can apply our procedure to find an approximately minimum volume ball on the transformed points, which will return a strong approximation since the variance of the points is controlled (via Lemma 2.1). Transforming this ball back into the original space will result in a confidence region that is shaped as an ellipsoid.
@@
    $\widehat{M}\succcurlyeq I$, and therefore $\widehat{M}^{-1/2}$ is “shrinking” (non-expanding) in all directions. This ensures that there exists a ball in the transformed space that contains $Y^{\star}$ and has volume at most that of $B^{\star}$ in the original space.
@@
Properties (i) and (ii) allow us to reduce to the non-worst-case setting, and property (iii) ensures that this translation does not lose too much in volume. We will set $\widehat{\tau}=d^{1/4}$ to trade-off the different losses from (ii) and (iii) and obtain the final guarantee. The full proof of this theorem can be found in Section 4.
We remark that our main result (Theorem 1.1) gives a volume approximation guarantee of $\exp(d^{1/2+o(1)})$ in the worst-case. At the same time, Corollary 1.4 gives a $O(1)$ factor approximation guarantee in volume, when the distribution of points inside the optimal ball $B^{\star}$ is approximately isotropic. We leave as an open question whether or not there exists a polynomial time (improper) learning algorithm that obtains an $O(1)$ volume approximation in the worst-case.
@@ 3 Proper Learning of Euclidean Balls
We use this result as a subroutine in our main result for improper learning. In this section we provide some additional results that extend our result for proper learning to other settings. In particular, Theorem 3.1 gives a proper learning algorithm for fully worst-case instances, and achieves an approximation ratio of $\left(1+\widetilde{O}_{\gamma,\delta}\left(\frac{1}{\log d}\right)\right)$ in radius, which is comparable to the factor of $\left(1+\widetilde{O}_{\gamma,\delta}\left(\frac{1}{\log d}\right)\right)$ due to (Badoiu et al., 2002; Ding, 2020). While we do achieve a slightly weaker approximation guarantee in this setting, our techniques are significantly different than the coreset based techniques of the prior work, and extend well to non-worst-case settings, as evidenced by our improper learning result. We also provide Theorem 1.4, which shows a way to achieve a better approximation ratio, assuming non-worst-case properties only on the inlier points.
@@
$\operatorname*{\mathrm{vol}}(\widehat{B})^{1/d}\leq\operatorname*{\mathrm{vol}}(B^{\star})^{1/d}\left(1+O\left(\frac{\log\log d}{\gamma\delta\log d}\right)\right)$
@@
$\operatorname*{\mathrm{vol}}(\widehat{B})^{1/d}\leq\operatorname*{\mathrm{vol}}(B^{\star})^{1/d}\left(1+O\left(\frac{\log\log d}{\gamma\delta\log d}\right)\right)$
@@
Our approach starts with the following structural observation about the true high-density ball $B^{\star}$. It is not in general true that the mean $\mu^{\star}$ of the points $Y^{\star}=B^{\star}\cap Y$ should be near the center $c^{\star}$ of $B^{\star}$. However, it turns out that if the mean is far from the center, taking a ball centered at the mean of $B^{\star}$ must still capture most of the points in $Y^{\star}$. This is a simple consequence of Chebyshev’s inequality— if we are able to control the variance of the points in $Y^{\star}$ (Lemma 2.1, Figure 1). Thus our problem reduces to the following. Given only the full set of points $Y$, we must estimate the mean of $Y^{\star}$, while also controlling the variance of the points that we are taking the mean of.
@@
Suppose there exists a subset of points $Y^{\star}\subseteq Y^{\prime}$ are contained in a smaller ball $B^{\star}=B(c^{\star},R^{\star})$, where $|Y^{\star}|\geq\delta|Y^{\prime}|$ and $R^{\prime}\leq 2R^{\star}$. Let $\gamma>0$ be a coverage slack factor.
Then there exists a list $C$ of size at most $\left(\frac{3R^{\prime}}{\tau}\right)^{q}$ and a set $\widehat{Y}\subseteq Y^{\star}$ with $|\widehat{Y}|\geq(1-\gamma)|Y^{\star}|$, such that there exists a point $\widehat{c}\in C$ which for every $y\in\widehat{Y}$ satisfies
$||y-\widehat{c}||_{2}^{2}$ $\leq||\Pi_{\mathrm{low}}(y-c^{\star})||_{2}^{2}+\frac{\sigma^{2}}{\gamma}\left(1+\frac{2(1-\delta)}{\delta}\right)+\left(\|\Pi_{\mathrm{high}}(y-c^{\star})\|_{2}+\tau\right)^{2}$
$\leq\left(||y-c^{\star}||_{2}+O\left(\frac{\sigma}{\sqrt{\gamma\delta}}\right)+\tau\right)^{2},$
where $\Pi_{\mathrm{high}}$ is the projection onto the top-$q$ eigenspace of $\Sigma_{Y^{\prime}}$ and $\Pi_{\mathrm{low}}$ is the projection onto the space orthogonal to the top-$q$ eigenspace of $\Sigma_{Y^{\prime}}$. For $q=0$, we get a single estimate $\widehat{c}$, and $\widehat{Y}\subseteq Y^{\star}$ with $|\widehat{Y}|\geq(1-\gamma)|Y^{\star}|$, that for every $y\in\widehat{Y}$ satisfies
$||y-\widehat{c}||_{2}^{2}$ $\leq||y-c^{\star}||_{2}^{2}+\frac{\sigma^{2}}{\gamma}\left(1+\frac{2(1-\delta)}{\delta}\right).$
@@
We start by considering $\Sigma_{Y^{\prime}}$, the covariance matrix of $Y^{\prime}$ and partition its eigenvalues into two groups. The directions associated with the largest $q$ eigenvalues of $\Sigma_{Y^{\prime}}$ are referred to as the “high variance directions”; while the directions corresponding to the remaining eigenvalues are the “low variance directions”. Let $\Pi_{\mathrm{high}}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{q}$ denote the projection onto the “high variance” eigenspace and $\Pi_{\mathrm{low}}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d-q}$ denote the projection onto the “low variance” eigenspace. We estimate the mean $\mu^{\star}$ of $Y^{\star}$ separately in the high variance and low variance directions.
First, if $q>0$, we estimate the location of $\mu^{\star}$ in the high variance directions. Let $Y^{\prime}_{\mathrm{high}}=\Pi_{\mathrm{high}}Y^{\prime}$ denote the projection of points in $Y^{\prime}$ onto the high variance directions. Note that the set $Y^{\prime}_{\mathrm{high}}$ lies within the $q$-dimensional ball $B^{\prime}_{\mathrm{high}}=\Pi_{\mathrm{high}}B^{\prime}$, which is centered at the origin and has a radius of at most $R^{\prime}$. Next, consider the projection of $Y^{*}$ onto the high variance directions, denoted by $Y^{*}_{\mathrm{high}}=\Pi_{\mathrm{high}}Y^{*}$. Then, this set $Y^{*}_{\mathrm{high}}$ is contained in the ball $B^{*}_{\mathrm{high}}=\Pi_{\mathrm{high}}B^{*}$ centered at $\Pi_{\mathrm{high}}c^{*}$ with radius at most $R^{\star}$. Since $B^{\star}_{\mathrm{high}}\subseteq B^{\prime}_{\mathrm{high}}$, and $\mu^{\star}\in B^{\prime}_{\mathrm{high}}$ as well, by Lemma 3.5, there exists a list $L\subset\mathbb{R}^{q}$ with size $|L|\leq\left(\frac{2+\tau/R^{\prime}}{\tau/R^{\prime}}\right)^{q}\leq\left(\frac{3R^{\prime}}{\tau}\right)^{q}$ such that there is a point ^$\widehat{\ell}\in L$ such that for every point $y\in Y^{\star}$,